Walt made an extra $9000 last year from a part-time job. He invested part of the money at a 9% annual rate of interest, and the rest at an 8% annual rate of interest. At the end of the year, he made a total of $770 in interest. How much was invested at 8%?
Answer by jorel1380(3719)

m+n=9000
.09m+.08n=770
9m+8n=77000
8m+8n=72000
m=5000
n=4000
$4000 was invested at 8%..